INAF LBC+LUCIFER queue observing, 2010 January 16-17

Observer: V. Testa, F. Mannucci
Telescope Operator: WWack
Instrument: LBC

Summary

* Thin and thick clouds all the night, seeing around 1"
* ~6 hours of scientific data, partly affect by low transparency
* Time lost: 30 min at UT=2 for collimation, 40min for clouds, 15min for "fiber link failure" at UT=13

Data Summary

Project Objects Filter ExpTime Notes
Palazzi GRB070714 U 15min problems with collimation
Palazzi GRB070714 V 30min  
Palazzi GRB070714 B 30min  
Palazzi GRB070714 U 30min  
Giallongo Q0933 G 4.0 h  
Giallongo COSMOS G 40 min  

Details

01:10 - clouds everywhere
01:20 - opening dome, LBC cannot read ALT/AZ from ETC
01:55 - much better sky, going to a pointing star; seeing~1.1"
02:05 - GRB070714 - Palazzi, filtro U,
02:20 - problems with collimation from the beginning, the guide star eppear double or out of focus. Integration aborted after ~15min, new collimation of the telescope
02:25 - GRB070714, filter B
02:37 - same problems with collimation , lost focusing, aborted
02:40 - GRB070714, filter V, collimation is conserved for ~20 min, seeing~1.3"
03:25 - GRB070714, filter U, 30 min, seeing~1", collimation OK
04:10 - GRB070714, filter B, 30 min, seeing~1"
05:00 - STD: SA95, UUsBVgr, AM=1.3
05:15 - pointing and collimation
05:30 - Q0933 - Giallongo, G, field=A AM=1.5, seeing~1.0", thin clouds
06:05 - collimation. Clouds (~3mag of attenuation), running the OB to monitor transparency
06:50 - better transparency, starting again
06:50 - Q0933, G, 30 min, field B, seeing~0.9"
07:40 - Q0933, G, 30 min, field C, seeing~1.2"
08:20 - Q0933, G, 30 min, field D, seeing~1.0", thin clouds
09:00 - Q0933, G, 30 min, field E, seeing~0.9", thicker clouds passing
09:40 - Q0933, G, 30 min, field F
10:30 - pointing: no stars because of clouds 10:45 - Q0933, G, 30 min field G
11:20 - Q0933, G, 30 min, B again to compensate for lower transmission
12:00 - COSMOS - Giallongo, G, field A, seeing=1.2"
12:40 - COSMOS - Giallongo, G, field B
13:00 - trackers stopped after 5 images, "fiber link failure"
13:15 - sky brightening, night ends

-- FilippoMannucci - 17 Jan 2010
Topic revision: r4 - 17 Jan 2010, FilippoMannucci
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback